OpenStudy (anonymous):

There were 190 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $530. How many Juniors bought tickets? 70 80 90 100

OpenStudy (anonymous):

Let the number of Juniors be x. So number of seniors = (190 - x) 2x + 3.5(190 - x) = 530 2x + 665 - 3.5x = 530 -1.5x = -135 Can you find x now? That will be the answer.

OpenStudy (anonymous):

112.50

OpenStudy (anonymous):

That's wrong! Think practically How can 112.5 people(Juniors) attend the dance?

OpenStudy (anonymous):

i dont know!! thats why i need help!

OpenStudy (anonymous):

Let x be the number of junior tickets Let y be the number of senior tickets 2x+3.5y=530----1 x+y=190 2x+2y=380------2 (1-2): 1.5y=530-380 =150 y=100 x=90 U can check if its correct

OpenStudy (anonymous):

@kiamonstaa You had divided -135 by -1.2 instead of -1.5. That's how you got 112.50.

OpenStudy (anonymous):

ohh haha i got it now!